# PAHA: Parts-Aware Audio-Driven Human Animation with Diffusion Model

### v4

@@ Abstract.
Audio-driven human animation technology is widely used in human-computer interaction, and the emergence of diffusion models has further advanced its development. Currently, most methods rely on multi-stage generation and intermediate representations, resulting in long inference time and issues with generation quality in specific foreground regions and audio-motion consistency. These shortcomings are primarily due to the lack of localized fine-grained supervised guidance. To address above challenges, we propose PAHA, an end-to-end audio-driven upper-body human animation framework with diffusion model. We introduce two key methods: Parts-Aware Re-weighting (PAR) and Parts Consistency Enhancement (PCE). PAR dynamically adjusts regional training loss weights based on pose confidence scores, effectively improving visual quality. PCE constructs and trains diffusion-based regional audio-visual classifiers to improve the consistency of motion and co-speech audio. Afterwards, we design two novel inference guidance methods for the foregoing classifiers, Sequential Guidance (SG) and Differential Guidance (DG), to balance efficiency and quality respectively. Additionally, we build CNAS, the first public Chinese News Anchor Speech dataset, to advance research and validation in this field. Extensive experimental results and user studies demonstrate that PAHA significantly outperforms existing methods in audio-motion alignment and video-related evaluations. The codes and CNAS dataset will be released upon acceptance.
@@
copyright: acmlicensedjournalyear: 2025conference: Proceedings of the 33nd ACM International Conference on Multimedia; October 27-31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33nd ACM International Conference on Multimedia (MM ’25), October 27-31, 2025, Dublin, Irelandsubmissionid: 1043ccs: Computing methodologies Computer vision
@@ 5. Experiments
Figure 6. A qualitative comparison highlights flaws in earlier methods, such as background inconsistencies, hand blurriness, and finger distortion (highlighted in red). In contrast, our method generates high-quality videos.
@@ 5.2. Quantitative Results
5.3. Qualitative Results
We qualitatively evaluate video generation in Fig. 6. The results indicate that SDT produces poor-quality faces with distorted expressions, while MM-Diffusion generates unrealistic shoulder movements. ANGIE’s gestures are inconsistent with the accompanying speech, and though S2G improves on ANGIE, it struggles with details like hands due to its reliance on unsupervised TPS keypoints for global transformation, neglecting local deformations and causing blurry outputs. In contrast, our method generates high-quality videos, especially in the face and hand regions, and significantly improves motion-speech consistency by utilizing a carefully trained PCE classifier for guidance during inference.
@@ Appendix A Overview
To better understand PAHA and clearly view the results, we provide a presentation_video.mp4 showcasing our core contributions (PAR, PCE, Inference Process) and visualizations.
@@ B.1. UniVDM
UniVDM is initialized from a pretrained video diffusion model (Blattmann et al., 2023). During training with the PAR method, videos have a spatial resolution of 512×512., a fixed length of 32 frames, a batch size of 8, and 100k training steps. We use the AdamW optimizer (Loshchilov, 2017) with a learning rate of 5e-5 and DDPM (Ho et al., 2020) as the noise scheduler with 1000 sampling steps. The confidence score threshold $\tau_{j}$ is set to 0.8, with a radius $r$ of 10. The hand/face weight hyperparameter $\omega_{1}$ is set to 10, while $\omega_{2}$ for the body region is set to 2.

### v5

# A Unit Enhancement and Guidance Framework for Audio-Driven Avatar Video Generation

@@ Abstract.
Audio-driven human animation technology is widely used in human-computer interaction, and the emergence of diffusion models has further advanced its development. Currently, most methods rely on multi-stage generation and intermediate representations, resulting in long inference time and issues with generation quality in specific foreground regions and audio-motion consistency. These shortcomings are primarily due to the lack of localized fine-grained supervised guidance. To address above challenges, we propose Parts-aware Audio-driven Human Animation, PAHA, a unit enhancement and guidance framework for audio-driven upper-body animation. We introduce two key methods: Parts-Aware Re-weighting (PAR) and Parts Consistency Enhancement (PCE). PAR dynamically adjusts regional training loss weights based on pose confidence scores, effectively improving visual quality. PCE constructs and trains diffusion-based regional audio-visual classifiers to improve the consistency of motion and co-speech audio. Afterwards, we design two novel inference guidance methods for the foregoing classifiers, Sequential Guidance (SG) and Differential Guidance (DG), to balance efficiency and quality respectively. Additionally, we build CNAS, the first public Chinese News Anchor Speech dataset, to advance research and validation in this field. Extensive experimental results and user studies demonstrate that PAHA significantly outperforms existing methods in audio-motion alignment and video-related evaluations. The codes and CNAS dataset will be released upon acceptance.
@@ B.1. UniVDM
UniVDM is initialized from a pretrained video diffusion model (Blattmann et al., 2023). During training with the PAR method, videos have a spatial resolution of 512×512, a fixed length of 32 frames, a batch size of 8, and 100k training steps. We use the AdamW optimizer (Loshchilov, 2017) with a learning rate of 5e-5 and DDPM (Ho et al., 2020) as the noise scheduler with 1000 sampling steps. The confidence score threshold $\tau_{j}$ is set to 0.8, with a radius $r$ of 10. The hand/face weight hyperparameter $\omega_{1}$ is set to 10, while $\omega_{2}$ for the body region is set to 2.